OpenStudy (anonymous):

A lighthouse is located on a small island 3km away from the nearest point P on a straight shoreline and its light makes 4 revolutions per minute. How fast is the beam of light moving along the shoreline when it is 1 km from P?

OpenStudy (anonymous):

.6 km per minute

OpenStudy (anonymous):

hello @TranceNova

OpenStudy (anonymous):

\[\tan \theta=y/x\]

OpenStudy (anonymous):

\[\sec^2\theta(\frac{ d \theta }{ dt }=\frac{ xdy }{ dt }-\frac{ ydx }{ dt } /(x^2)\]

OpenStudy (anonymous):

the right side is all over x^2

OpenStudy (anonymous):

\[\frac{ dy }{ dt }=\frac{ x^2(\sec^2\theta)\frac{ d \theta }{ dt } +y\frac{ dx }{ dt }}{ x }\]

OpenStudy (trancenova):

yep yep

OpenStudy (anonymous):

\[\frac{ 3^2(\frac{ 1 }{ \cos^2\theta })(8\pi)+1(0) }{ 3 }\]

OpenStudy (anonymous):



OpenStudy (anonymous):

@TranceNova my question to you is, whats the answer for 1/cos^2o?

OpenStudy (anonymous):

my book says the answer is 8pi/3 km/min but isnt cos^2theta (3/root10)^2?

OpenStudy (anonymous):

@bahrom7893 hey

OpenStudy (anonymous):

so im pretty much at the end, and I want to now what i am messing up :) thanks for listening

OpenStudy (bahrom7893):

Wait I'm kinda lost.. let me reread the problem.

OpenStudy (bahrom7893):

im getting 80pi/3

OpenStudy (anonymous):

so it is 10 right?

OpenStudy (bahrom7893):

http://answers.yahoo.com/question/index?qid=20110308160630AAZ3vZc This answer summarizes it nicely

OpenStudy (anonymous):

ok good, thank you so much @bahrom7893 !